OpenStudy (anonymous):

What is the solution to this sytem of equations: y − 2x = −1 x + 3y = 4

OpenStudy (shane_b):

Do you have any clue where to start?

OpenStudy (anonymous):

I changed it to slope intercept but the y intercept came out to 4 over 3?

OpenStudy (shane_b):

I would simply use substitution to solve this: Looking at the first equation: y-2x=-1 Which can be easily rearranged in terms of y: y=2x-1 Plug this equation into the second original equation to get: x + 3(2x-1) = 4 Solving for x you will get 1. Now just plug that into one of the original equations and solve for y: y − 2(1) = −1 y = 1.